OpenStudy (anonymous):

16p - 4 = 4 (2p - 3) ((show work))

OpenStudy (anonymous):

new look 16p-4=0.5

OpenStudy (anonymous):

0.5+4=4.5 subtract opposite is add

OpenStudy (anonymous):

16/16=1 no need for this but we do need is 4.5/16=0.28 p=0.28

OpenStudy (anonymous):

Expand the right side: 16p-4=8p-12 subtract 8p from both sides and add 4 on both sides to isolate p: 8p=-8 p = -1

OpenStudy (anonymous):

but u need to opposite of multiply

OpenStudy (anonymous):

don't u

OpenStudy (anonymous):

@ozzy22 dude... you don't need to do opposites for everything. this is basic algebra.